# Critical re-examination of a claimed challenge of Bohmian mechanics

### v3

@@ Abstract
A recent experiment by Sharoglazova et al. was interpreted by the authors as a challenge to the basic idea of Bohmian mechanics, the phase-velocity relation $\vec{v}(\vec{r},t)=\frac{\hbar}{m}\vec{\nabla}S(\vec{r},t)$. Here, we reinterpret their experiment, highlighting the role of the transient regime before the onset of the stationary state. In this way, we show that the experimental data can naturally fit the Bohmian framework when the energy in the evanescent regime is identified with the quantum potential. It is, however, also possible to interpret the same energy as a kinetic energy, in the framework of Nelson’s stochastic quantum mechanics. In conclusion, far from challenging the Bohmian mechanics, we find that the experimental results of Sharoglazova et al. can be interpreted in either ontologies and are therefore not conclusive in selecting one or the other.
@@ I Introduction
Recently, Sharoglazova et al. [1] claimed to have proven a violation of $\vec{v}(\vec{r},t)=\frac{\hbar}{m}\vec{\nabla}S(\vec{r},t)$, the phase-velocity relation of Bohmian mechanics [2], by measuring the speed [3] of an evanescent wave, which is real and therefore must have $\vec{\nabla}S(\vec{r},t)=\vec{0}$ (here, $S(\vec{r},t)$ is the real phase of the wave function $\psi(\vec{r},t)=\rho^{\frac{1}{2}}(\vec{r},t)e^{iS(\vec{r},t)}$).
The main idea at the basis of this measurement is explained in [4]: consider the composed waveguide used in the experiment, schematically depicted in Fig. 1. In the evanescent regime ($E<V_{0}$, with $V_{0}$ the potential step for $x\geq 0$), the wave in $w_{1}$ from $x<0$ penetrates the step at $x=0$ around $y=+a$ as an exponentially decaying wave for $x>0$ (see also Fig. 2). At the same time, it is experimentally found to fill in also the second waveguide ($w_{2}$) with a profile growing with $x$. This, naturally leads to the question: if some density enters the $x>0$ region in the first waveguide without any apparent velocity (in the evanescent state $\rightarrow\vec{\nabla}S=\vec{0}$), and then a density profile is measured in the second waveguide, indicating a spectral-wave displacement from $w_{1}$ to $w_{2}$, can we conclude that there is a velocity associated to this process? and how to combine this velocity to $\vec{\nabla}S=\vec{0}$? Interestingly, [1, 4] have also been able to quantify the population-transfer dynamics through a well-defined expression for the speed (see Section III), giving, in their words [4], ’a clear physical meaning to the notion of a speed for classically forbidden motion’.

### v4

# Critical re-examination of a claimed challenge to Bohmian mechanics

@@ Abstract
We re-analyze a recent experiment by Sharoglazova et al. highlighting the role of the transient regime. We prove that in the evanescent state of the stationary regime their experimental data can be interpreted in terms of Bohmian quantum mechanics, at odds with the challenge posed. At the same time, Bohm’s quantum potential can be re-interpreted as a kinetic-energy term in the framework of Nelson’s stochastic quantum mechanics, with a hidden-variable, non-classical, speed fitting the experimental data as well. The experiment can be interpreted as well within orthodox quantum mechanics and is therefore not conclusive in selecting or challenging any framework.
@@ I Introduction
Recently, Sharoglazova et al. [1] claimed to have proven a violation of the phase-velocity relation of Bohmian mechanics [2], $\vec{v}(\vec{r},t)=\frac{\vec{\nabla}S(\vec{r},t)}{m}$, by reporting the measurement of a finite speed for an evanescent wave. The latter being real it implies $\vec{\nabla}S(\vec{r},t)=\vec{0}$, where $S(\vec{r},t)$ is the phase of the wave function $\psi(\vec{r},t)=\rho^{1/2}(\vec{r},t)e^{iS(\vec{r},t)/\hbar}$.
The basic idea behind their claim is explained in [3]: the authors realized a waveguide composed of several parts, schematically depicted in Fig. 1. In the challenging part of their experiment a propagating wave with energy $E<V_{0}$ enters the waveguide $w_{1}$ located around $y=+a$ from $x<0$. Then, it encounters the potential step $V_{0}$ existing for $x\geq 0$. The exponentially decaying wave experimentally measured in $w_{1}$ at $x>0$ is also found to fill in the second waveguide $w_{2}$ located around y=ay=-a. The density profile in $w_{2}$ grows with $x$ for small $x$ (see also Fig. 3). So, the reported measurements seem to imply that at $x\geq 0$ some density moves across the waveguides, from $w_{1}$ to $w_{2}$. The latter would result at odds with the phase-velocity relation of Bohmian mechanics for a real evanescent wave. This naturally leads to the challenge proposed by Sharoglazova et al., who also quantified in terms of a speed the measured stationary density profiles. In their own words, the key hypothesis reads [1]: ’If the probability amplitude is initially concentrated entirely in one of the states [here, $w_{1}$], the population in the initially unoccupied state [here, $w_{2}$] follows $\sin^{2}(J_{0}t)$ as a function of time’, from which the deduction of the $w_{1}$ to $w_{2}$ motion at any time t.